### v1

Abstract
Identifying procedural errors online from egocentric videos is a critical yet challenging task across various domains, including manufacturing, healthcare, and skill-based training. The nature of such mistakes is inherently open-set, as unforeseen or novel errors may occur, necessitating robust detection systems that do not rely on prior examples of failure. Currently, however, no technique effectively detects open-set procedural mistakes online.
We propose a dual-branch architecture to address this problem in an online fashion: one branch continuously performs step recognition from the input egocentric video, while the other anticipates future steps based on the recognition module’s output. Mistakes are detected as mismatches between the currently recognized action and the action predicted by the anticipation module. The recognition branch takes input frames, predicts the current action, and aggregates frame-level results into action tokens. The anticipation branch, specifically, leverages the solid pattern-matching capabilities of Large Language Models (LLMs) to predict action tokens based on previously predicted ones.
Given the online nature of the task, we also thoroughly benchmark the difficulties associated with per-frame evaluations, particularly the need for accurate and timely predictions in dynamic online scenarios.
Extensive experiments on two procedural datasets demonstrate the challenges and opportunities of leveraging a dual-branch architecture for mistake detection, showcasing the effectiveness of our proposed approach. In a thorough evaluation—including recognition and anticipation variants and state-of-the-art models—our method reveals its robustness and effectiveness in online applications.
keywords:
Egocentric Vision, Procedural Mistake Detection, Video Understanding, Large Language Models, Chain of Thought, In-Context Learning
1 Introduction
Detecting procedural errors in videos has recently gained significant attention due to its potential to provide substantial benefits across various domains. Accurate mistake detection in videos of individuals performing tasks—such as recipe execution, object assembly, or managing complex workflows—holds promise for enhancing training and learning experiences by providing real-time feedback (see Figure 1). This immediate feedback allows for timely corrections, fostering faster skill development, skill acquisition, and a safer learning environment in high-stakes areas like surgery and aviation.
Figure 1: Procedural Mistake Detection involves identifying errors within a procedural video, where each procedure consists of steps that can be correct or incorrect. The aim is to develop a method capable of analyzing a video and determining whether each frame contains a mistake. This concept is particularly useful when applied to egocentric videos, as they allow for direct feedback to the person performing the task. In the diagram, the leftmost column shows the steps that define a procedure, while the right side presents a sequence of frames, which can either indicate a mistake (red) or a correct action (green). The black arrow illustrates the flow of time, and the algorithm, depicted by a gear icon, takes the video sequence from time 00 to $t-1$ as input, classifying the current frame $t$ as either correct (green) or a mistake (red).
Advanced mistake detection models will become essential for ensuring precision, safety, and efficiency in procedural applications. This demand has driven the creation of new datasets [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] and methodologies designed to advance procedure learning [13, 4, 14, 15, 1, 6, 16] and error detection models [3, 5, 8]. Currently, existing mistake detection approaches vary widely. Some methods emphasize action detection, identifying specific errors like missing steps or out-of-sequence actions [17], or incorrect oredering [16]. While other models bypass action detection altogether and monitor changes to the assembled object to ensure procedural accuracy [18, 19].
An ideal Procedural Mistake Detection (PMD) model should demonstrate two critical capabilities: robustness to diverse mistake types and the ability to provide online, timely feedback. First, a mistake detection system should identify any deviation from the correct procedure, regardless of the error’s nature. Since the range of correct methods is finite while possible mistakes are virtually unlimited, we use the One-Class Classification (OCC) framework. OCC trains on correct procedure videos only and tests on a mix of correct and incorrect cases, similar to anomaly detection [20, 21]. Second, the model should provide online feedback quickly enough to allow users to correct mistakes promptly, minimizing the reinforcement of incorrect actions.  [^1] We distinguish an online setup, which allows slight processing delays while enabling effective feedback, from a real-time setup, which requires immediate responses. Additionally, an egocentric perspective enhances the model’s applicability by mirroring natural human perception, thereby increasing feasibility for real-world deployment in task execution scenarios.
We propose a dual-branch approach combining step recognition and anticipation to detect mismatches in predicted and observed actions. The first branch, a step recognition module, processes the input video to identify the current action being performed. The second branch, a step anticipation module, predicts the next step based on previously recognized actions, utilizing a language model’s reasoning capabilities. Mistakes are flagged when there is a divergence between recognized and anticipated actions, highlighting inconsistencies between actual and expected steps.
This paper extends our previous conference work, PREGO [22], by refining key components to improve procedural mistake detection. While PREGO’s evaluation of Large Language Models (LLMs) focused on a single model and fixed approach, our work conducts a comprehensive evaluation of LLMs as step anticipators, achieving state-of-the-art results on Assembly101-O and Epic-tent-O [22]. Specifically, we examine a broader range of LLMs, fine-tune them using Low-Rank Adaptation (LLoRA), and experiment with various prompting strategies, including Zero-Shot, Few-Shot, and Automatic Chain of Thought (Auto-CoT) [23], with Auto-CoT significantly enhancing the anticipation branch’s predictive abilities.
Further, we address noise in the recognizer’s frame-by-frame predictions, which can disrupt action sequence continuity. To mitigate this, we extend beyond fixed-window aggregation to explore alternative strategies that reduce noise and improve prediction accuracy. We also introduce frame-level metrics for more precise evaluation, aligning the recognition module’s output with the anticipation module’s input. Given the online nature of this task, we emphasize accurate, timely per-frame evaluations in dynamic scenarios to ensure effective mistake detection and reduce response delays.
Our contributions include (1) exploring and benchmarking multiple LLMs to advance PREGO’s foundational approach, (2) experimenting with various prompting techniques, including Auto-CoT, which enhances overall performance, (3) evaluating frame-level metrics for more precise system assessment compared to the original fixed window approach, and (4) investigating alternative aggregation techniques to improve mistake detection robustness and continuity. Extensive experimentation demonstrates the challenges and opportunities in dual-branch architectures and LLM-based step anticipation for open-set mistake detection. Our results underscore the potential of integrating step recognition and anticipation within a unified framework for online procedural error detection, offering a more flexible and effective solution for real-world applications.
2 Related Works
2.1 Procedural Mistake Detection
Procedural learning has made notable progress with the development of diverse datasets [6, 10, 9, 11, 12], offering valuable insights into both structured [7, 3, 24] and unstructured [1, 25] tasks. These datasets span a wide range of applications, from industrial assembly [24, 2, 8, 12] to everyday cooking activities [25, 26, 27]. Despite the growing interest in this field, the lack of a standardized method for mistake detection has resulted in limited evaluation and fragmented literature.
2.2 Datasets of Procedural Egocentric Videos
Assembly101 [2] is a large-scale video dataset that provides frame-level mistake annotations. It features videos of actors assembling toys, with synchronized Ego-Exo views and annotated hand positions. IndustReal [8] focuses on a single toy, which leads to a single procedure being learned. Epic-tent [1] is a dataset that covers a different domain of unscripted actions that capture actors assembling a tent outdoors. The participants exhibit varying levels of expertise and naturally make mistakes, which have been annotated. ATA [4] is a procedural dataset created for offline mistake detection in assembling activities. However, it provides only video-level mistake annotations, limiting its usefulness for frame-based applications. HoloAssist [3] provides egocentric videos of individuals performing multiple manipulation tasks following expert instructions. CaptainCook4D [28] is tailored for evaluating procedural activities, specifically focusing on the culinary domain. While comprehensive, [28] incorporates both procedural errors and attentional lapses. For our purposes, the latter category falls outside the scope of our investigation, yielding such dataset unsuitable. Notably,  [29] introduced a novel egocentric procedural error dataset consisting of videos depicting various errors within the cooking domain. In this study, we propose a new benchmark building upon PREGO tilize the datasets from [2, 1] as they provide insights into procedural errors in two distinct settings: controlled industrial and outdoor environments.
Table 1: Comparison of relevant models in procedural mistake detection. In the modalities column, RGB refers to RGB images, H stands for hand poses, E represents eye gaze, and K indicates keystep labels. Our approach is the first to adopt an egocentric, one-class and online method for detecting procedural mistakes. (*) denotes that these models are based on our protocol defined in [22].
Model Egocentric (Ego) One-Class (OCC) Online Modalities Task Datasets
Ding et al. [5] - ArXiv ’23 K Mistake Detection Assembly101 [2]
Wang et al. [3] - ICCV ’23 RGB+H+E Mistake Detection HoloAssist [3]
Ghoddoosian et al. [4] - ICCV ’23 RGB Unknown Sequence Detection ATA [4], CSV [30]
Schoonbeek et al. [8] - WACV ’24 Multi Procedure Step Recognition IndustReal [8]
Lee et al. [29] - CVPR ’24 RGB Mistake Detection EgoPER [29], HoloAssist [3], ATA [4]
Seminara et al. [16] - NeurIPS’24 * RGB Mistake Detection Assembly101-O, Epic-tent-O
PREGO - CVPR ’24 * [22] RGB Mistake Detection Assembly101-O, Epic-tent-O
TI-PREGO * RGB Mistake Detection Assembly101-O, Epic-tent-O
2.3 Methods for procedural mistake detection
Table 1 presents an overview of the core characteristics of recent methods for Mistake Detection in procedural videos. Ding et al. [5] propose a method for procedural learning that utilizes knowledge graphs to identify errors. Their approach is based on textual transcripts, ignoring the use of visual cues coming from videos.
In [4], the authors train an action recognition model and treat error detection as a semantic evaluation of the segmentation results. On the other hand, Assembly101 [2] and HoloAssist [3] apply the same error detection baselines on varying granularity but also run offline, requiring video segmentation and labels during training, thereby limiting their ability to detect previously unseen errors. In  [8], the authors address the problem of Procedure Step Recognition, which focuses on recognizing the correct completion of steps rather than the partial execution of activities. EgoPER [29] proposes a framework composed of an action segmentation module and a constrastive module for detecting errors unseen during training. PREGO [22] leverages video frames to detect procedural steps and incorporates symbolic reasoning for evaluating the procedure’s correctness. Its objective is to identify mistakes online as soon as they occur, ensuring prompt detection. Additionally, it is designed to function within an OCC framework, which allows it to detect a wide range of errors since it is trained on sequences without mistakes, without any constraint on specific errors  [20, 21].
Seminara et al. [16] learn task graphs by calculating the probability of witnessing a sequence of key steps. In contrast to the proposed methodology, which depends solely on context without further training for the anticipator, they employ a training approach grounded in Maximum Likelihood estimation.
2.4 Step recognition
Step recognition, the process of identifying discrete actions within a structured procedural sequence, is fundamental to progress in domains such as autonomous robotics and educational technology.
Zhuang et al. [13] introduce an action segmentation model that employs an attention-based architecture coupled with a Pairwise Ordering Consistency (POC) loss function. This approach effectively captures the correct sequence of procedural steps. Their innovation lies in developing a weakly-supervised method that requires only the set of actions in a procedure as input, eliminating the need for labor-intensive frame-level annotations. In Shah et al.[31], a novel loss for self-supervised learning is combined with a clustering algorithm to detect key steps in procedural videos without using any labels. [15] tackles the task by utilizing online instructional videos to learn actions and sequences without manual annotations, combining step recognition with a deep probabilistic model to account for step order and timing variability. An et al. [32] introduced MiniROAD, the state-of-the-art online step recognizer, which utilizes an RNN architecture and adjusts loss importance during training to perform online action recognition. However, existing methods like MiniROAD perform detection on a frame-by-frame basis, which can be limiting in practical scenarios that require recognizing sequences of actions rather than isolated frames. This limitation is particularly apparent in real-world situations where understanding the broader context of sequential actions is crucial. To overcome this, we introduce aggregation techniques to improve the coherence of action recognition, addressing the frame-level limitation inherent in current methods.
Moreover, the approach proposed in Shen et al. [33] also highlights the importance of holistically understanding actions rather than treating each frame in isolation. Yet, while PREGO focuses on detecting procedural mistakes using an OCC-based approach with LLMs for action anticipation, [33] targets action segmentation and progress estimation with task graphs.
2.5 Step Anticipation
Conversely, step anticipation predicts upcoming actions in a sequence, which is important for real-time AI decision-making. [34] generates multiple possible natural language outcomes, leveraging pretraining on a text corpus to handle the variability in future actions. [35] employs a two-level hierarchical approach, integrating both high-level human intentions and low-level action sequences, improving action anticipation in the long term. The pipeline consists of two models: the first extracts intentions and classifies actions, and the second generates future actions, conditioning human intentions to narrow down the uncertainty set of future actions.
Furthermore, the framework presented in [36] addresses future activity anticipation in egocentric videos by employing a contrastive loss to emphasize novel information and a dynamic reweighting mechanism that prioritizes informative past content. This approach improves video representation, leading to more accurate predictions of future activities. While we adopt a similar dual-branch architecture as in PREGO  [22]. However, we omit symbolic reasoning, instead leveraging the actual semantic labels to harness the reasoning capabilities of LLMs through an Automatic Chain of Thought.
2.6 Reasoning tasks with Large Language Models
Large Language Models (LLMs), trained on vast datasets and equipped with numerous parameters, exhibit advanced capabilities beyond those of earlier language models [37]. They have demonstrated exceptional performance in both natural language processing tasks [38] and non-language tasks [39, 37, 40]. The next-token prediction mechanism of LLMs closely parallels our action anticipation framework, as both aim to predict future outcomes based on accumulated data.
Recent research [41, 40, 42, 43, 44, 45, 46] has investigated the ability of LLMs to function as In-Context Learners (ICLs), meaning they can solve novel and unseen tasks without requiring additional fine-tuning. When provided with a query prompt that includes a context of input-output examples, LLMs can understand the problem and generate appropriate solutions within this framework. LLMs as ICLs have been applied to a wide range of tasks, including planning [41, 46], programming [40, 42], logical problem-solving [43], and symbolic reasoning [44]. In addition to ICL, other paradigms, such as Chain-of-Thought (CoT) [47] and Automatic Chain-of-Thought (ACoT) [23], have been employed to enable LLMs to reason through their responses step-by-step, improving their ability to plan and articulate their thought process explicitly.
Previous research has shown that LLMs can generate semantically meaningful patterns [44, 48], while other work [49] has explored their in-context learning abilities with semantically unrelated labels, where there is no direct relationship between a token and its meaning. Recent studies [41, 50, 45] have investigated using LLMs for creating task-oriented plans
The approach in [45] is the most similar to ours, as it employs a Socratic method [51] that uses video features to predict subsequent actions with an LLM. Socratic Models (SMs) are a framework that utilizes structured dialogue between pre-existing foundation models, each leveraging its unique capabilities based on its training data distribution. However, while PALM integrates an Action Recognition Model (ARM) and a Vision-Language Model (VLM) to generate text prompts for long-term action anticipation, we diverge by using a dual-branch architecture: one branch dedicated to video recognition for action detection and the other utilizing an LLM for action anticipation.
In our mistake detection pipeline, we integrate In-Context Learning (ICL) and Automatic Chain of Thought (ACoT), utilizing an LLM as part of our action anticipation branch. ACoT enables the model to automatically generate intermediate reasoning steps, breaking the problem-solving process into smaller, logical steps. The approach helps the LLM to externalize its reasoning, enhancing its effectiveness in managing complex, sequential tasks.
3 Methodology
The proposed system leverages a dual-branch framework that integrates the recognition of procedural steps with anticipation modeling, as shown in Fig. 2. In the following sections, we elaborate on the problem formalization (Sec. 3.1), present the branches for step-recognition (Sec. 3.2) and step-anticipation (Sec. 3.3) and finally illustrate the mistake detection procedure (Sec. 3.4).
Figure 2: Our proposed model is based on two main components: The recognition module (orange) processes the input video in an online fashion and predicts actions observed at each timestep; the anticipation module (blue) reasons symbolically via a Large Language Model, utilizing automatic Chain of Thought (ACoT) reasoning to predict the future action based on past action history and a brief context, such as instances of other action sequences. Mistakes are identified when the current action detected by the step recognition method differs from the one forecasted by the step anticipation module.
3.1 Background
Let us consider a finite collection of $N$ procedures $\{p_{i}\}_{i=1}^{N}$, where each procedure encodes a sequence of actions as $p_{i}=\{a_{k}\}_{k=1}^{K_{i}}$. Here, $K_{i}$ varies depending on the specific procedure, and each action $a_{k}$ belongs to the set $a_{k}\in\mathcal{A}=\{a|a\text{ is a possible action}\}$. Additionally, each procedure is represented by a set of videos $\{v_{i}\}_{i=1}^{N}$, which consist of frames $v_{i}=\{f_{\tau}\}_{\tau=1}^{M_{i}}$, with $M_{i}$ indicating the total number of frames in video $i$. Given a specific frame $f_{\tau}$ from video $v_{i}$, our task is two-fold: we (1) identify the action $a_{\tau}$ corresponding to the frame $f_{\tau}$ and (2) predict the action $a_{\tau}$ that will occur at time $\tau$, relying solely on past observations up to time $\tau-1$.
The step recognition task is handled by a module $\rho$, which takes the encoded frames of $v_{i}$ from the start up to frame $\tau$ as input and outputs an action $a^{\rho}_{\tau}$ (Sec. 3.2). Subsequently, we input all the actions $a^{\rho}_{1},...,a^{\rho}_{\tau-1}$ into module $\xi$, which is responsible for the anticipation task, allowing it to predict the next action in the sequence as $a^{\xi}_{\tau}$ (Sec. 3.3).
Finally, we compare $a^{\rho}_{\tau}$ with $a^{\xi}_{\tau}$ designating actions as mistaken when there is a misalignment between the outputs from the two branches (Sec. 3.4). For clarity, the rest of this section will focus on a single procedure $p$ associated with a video $v$.
3.2 Step Recognition
The step recognition module, referred to as $\rho$, takes a window $W=\{{f_{{\tau}-w},..,f_{\tau}}\}$ with fixed length $w$ of encoded frames from $v$ up to the current frame $\tau$ as input, and is tasked to recognize the last step performed $a^{\rho}_{\tau}$. In our approach, we use MiniRoad [32], the state-of-the-art model for online step detection, which is also computationally efficient in terms of GFlops and parameter count.
In this setup, the model makes predictions for the action $a_{\tau}$ by inspecting the frames ${f_{{\tau}-w},..,f_{\tau}}$.
$\rho({f_{{\tau}-w},..,f_{\tau}})=a^{\rho}_{\tau}$ (1)
This technique often results in redundant predictions, with the model frequently assigning the same action to consecutive frames. To address this issue, we explore methods for aggregating redundant and similar actions in Sec. 6.
During training, the loss for this step recognition module is calculated through a Cross Entropy Loss, comparing the actual action $a_{\tau}$ with the predicted action $a^{\rho}_{\tau}$.
3.3 Step Anticipation
We employ an LLM as our $\xi$ module for next-step prediction, feeding it with prompts from procedural video transcripts and exploiting its reasoning ability enhanced with the Automatic Chain of Thought (ACoT) [23] technique. Our framework operates without specific training or fine-tuning, leveraging only the reasoning from ACoT (see Fig. 2) and few-shot prompts. These prompts consist of two components: the contextual transcripts $C$ and the current sequence $s_{\tau}$. The former is derived from similar procedures (i.e., different procedures for the same toy) to provide the LLM with information about typical step sequences and their order. The latter, Sequence $s_{\tau}$, includes the current sequence of actions up to the current frame $f_{\tau}$, as they have been detected by our module $\rho$, i.e., the sequence $s_{\tau}$ up to frame $f_{\tau}$ is defined as $s_{\tau}=[a^{\rho}_{1},...,a^{\rho}_{\tau-1}]$, where $a^{\rho}_{i}$ represents the action detected by the module $\rho$ at frame $i$.
The prompting occurs in two stages. First, we employ an ACoT mechanism of $\xi$ as an intermediate step $\phi_{\tau}$. This ACoT process uses $C$, $s_{\tau}$, and an additional prompt to explicitly stimulate reasoning. Then, we use the output of $\phi_{\tau}$ along with $C$ and $s_{\tau}$ to predict the next most probable action. By breaking down the task into smaller logical steps, the LLM can leverage the contextual transcript $C$ and the current action sequence $s_{\tau}$ to generate intermediate reasoning steps, bridging the gap between observed and future actions.
3.4 Mistake Detection
Ultimately, we analyze the outputs of the two modules to identify procedural errors. Specifically, we classify steps as correct when the outputs of both modules match, while we label cases as errors when the outputs differ. That is:
$\begin{cases}a^{\rho}_{\tau}\neq a^{\xi}_{\tau}&\text{MISTAKE}\\ a^{\rho}_{\tau}=a^{\xi}_{\tau}&\text{CORRECT}\\ \end{cases}$ (2)
We compare two evaluation settings: aggregated and frame-level. In both cases, the input data to the recognizer is aggregated; this is done by compressing multiple window frames’ actions in a single one by considering the most frequent action and then merging adjacent actions. While in the aggregated case, we directly compare the anticipator’s output to that aggregated input, in the frame-level case, we have to expand it to the original dimensionality.
4 Benchmark Setup
This section describes the benchmark datasets and evaluation metrics used in our experiments. We introduce Assembly101-O and Epic-tent-O as adaptations of the original datasets [2, 1], detailing the selected labeling criteria for online benchmarking, a novel configuration of training and test splits to accommodate open-set procedural mistakes, and defining our proposed online metrics in Sec. 4.1.
4.0.1 Assembly101-O
Assembly101 [2] is a large-scale video dataset that facilitates procedural video understanding. The dataset contains 362 videos of assembly and disassembly tasks involving 101 types of toy vehicles, recorded from multiple static (8) and egocentric (4) cameras. It includes extensive annotations at various levels of granularity, with over 100K coarse and 1M fine-grained action segments and 18M 3D hand poses. The dataset supports diverse challenges, including action anticipation and segmentation, mistake detection, and 3D pose-based action recognition.
Assembly101 for Online and Open-Set Mistake Detection (Proposed) We introduce a novel split of the dataset [2] tailored explicitly for online, open-set mistake detection. Assembly101-O introduces two main modifications to [2] as we show in Figure 3: a new train/test split and revised procedure lengths. The new split assigns all correct procedures to the training set while reserving videos with mistakes for validation and testing. This adjustment enables models to learn sequences characteristic of correct procedures in a one-class classification (OCC) framework, avoiding bias from specific mistake types during training. By training exclusively on correct sequences, models adhere to the OCC protocol, treating all deviations as mistakes and enabling better balance in the validation and test sets for more comprehensive mistake detection evaluation.
Figure 3: The modification occurs in two steps: starting from Assembly101 (a), first, a new train/test split assigns all correct procedures to the training set, reserving videos with mistakes for validation and testing (b). Second, the lengths of procedures containing mistakes are adjusted by trimming them to the first mistake, preventing the creation of corrupted sequences (c). This setup enables models to learn correct sequences within a one-class classification (OCC) framework, treating any deviation as a mistake and ensuring a balanced validation and test set for effective mistake detection.
The second adjustment limits each procedure’s evaluation to the point where a mistake caused by an incorrect action dependency disrupts the workflow. Per the OCC protocol, models are trained to recognize the correct step sequences necessary for completing procedures. Evaluating beyond the point of procedural compromise creates discrepancies between the training and test set actions, hindering effective step recognition and anticipation. Furthermore, to align with real-world applications, this work focuses on egocentric videos, utilizing only a single egocentric view per procedure from the four available in [2].
4.0.2 Epic-tent-O
Epic-tent is a dataset of egocentric videos capturing the outdoor assembly of a camping tent. It was collected from 24 participants who wore two head-mounted cameras (GoPro and SMI eye tracker) while performing the task, resulting in 5.4 hours of recorded video. The dataset includes annotations for action labels, task errors, participant self-rated uncertainty, and gaze positions, capturing the variability and complexity of the task as participants interacted with non-rigid objects (such as the tent, guylines, instructions, and tent bag) and displayed varying proficiency and confidence levels.
Unlike Assembly101 [2], Epic-tent is designed explicitly for supervised error detection, with each recorded procedure containing at least one mistake, making it incompatible with the split procedure proposed for Assembly101-O. However, Epic-tent provides frame-level confidence scores from each participant, reflecting their self-assessed uncertainty during task performance. Therefore, we define a custom split strategy (detailed in Sec. C of the supplementary materials): videos from higher-confidence participants form the training set. In contrast, participants showing more significant uncertainty (and thus potentially prone to more errors) make up the test set. This strategy is particularly promising for real-world applications where accurately labeling erroneous frames may be challenging or training a mistake detector can start soon after recording without requiring full annotation completion. This partition results in 14 training videos and 15 test videos.
Epic-tent for Online and Open-Set Mistake Detection (Proposed) This section introduces a novel split for the Epic-tent dataset [1] tailored for open-set mistake detection. The dataset includes annotations for nine distinct mistake types, though only four of these categories—“order”, “omit”, “correction”, and “repeat”—are considered procedural mistakes, as they represent deviations that compromise the task’s procedural integrity. Other categories such as “slow”, “search”, “misuse”, “motor”, and “failure” are not procedural errors, as they do not disrupt the task’s progression.
The Epic-tent dataset exclusively comprises egocentric videos recorded with Go-Pro cameras, reinforcing the dataset’s suitability for open-scene, real-world applications. Videos in the test set are trimmed to the frame, marking the first procedural mistake, while videos representing correct procedures remain unaltered.
4.1 Metrics Definition
We work with a sequence of frames, each representing a procedure, where all actions are correct except for the last one, which is a mistake. A mismatch between the prediction of the step recognition and the ones of the step anticipation indicates the presence of a mistake, while no mismatch means the procedure is correct. To assess the performance of our procedural mistake detection model, we follow the evaluation proposed in PREGO: we use True Positives as a measure of the model correctly identifying errors and True Negatives as a measure of accurately labeling steps that are not errors. Thus, we rely on the Precision, Recall, and F1 score metrics to evaluate the performance of our model. These metrics offer valuable insights into the model’s capability to identify and classify mistakes within procedural sequences. Precision quantifies the accuracy when predicting mistakes, minimizing false positives. Recall assesses the model’s capability to retrieve all mistakes, reducing the number of false negatives. Finally, the F1 score is the harmonic mean of precision and recall, and it balances failures due to missing mistakes and reporting false alarms.
To establish a method-agnostic evaluation metric, we focus solely on the model’s ability to classify actions as correct or erroneous. While our approach incorporates an online action detection component, we do not directly assess its general step recognition capabilities—instead, our evaluation centers on the accuracy of distinguishing between correctly and incorrectly performed actions. This decision stems from the limitations of metrics that blend recognizer and anticipator module performances, which can lead to unfair comparisons with approaches employing alternative prediction strategies. Given the expected class imbalance in ego-centric mistake detection tasks and the critical importance of identifying errors, the mistakes are considered as mistakes as the positive class in our evaluation framework.
When evaluating frame-level metrics (see Sec. 7.3), we consider corner cases such as the best, worst, and random scenarios (see Sec. 7.1) for clarity. To ensure a fair comparison, we use balanced accuracy as the evaluation metric, the average of true positive and true negative rates. This is useful to mitigate the impact of dataset imbalance, which would otherwise lead to an unfair comparison.
Balanced $\text{Balanced Accuracy}=\frac{1}{2}\left(\frac{TP}{TP+FN}+\frac{TN}{TN+FP}\right).$ (3)
5 Large Language Models as Step Anticipators
This section comprehensively analyzes how an action anticipator operates using various techniques and methodologies. We begin by evaluating the performance of several large language models (LLMs) using in-context learning with few-shot examples (Sec. 5.1), which allows us to determine which model performs best in this setting. Once the most effective LLM is identified, we investigate different prompting strategies, starting from a system prompt’s impact, including zero-shot, few-shot, and Automatic-Chain-of-Though prompting (Sec 5.2). In advance, we explore the potential improvements by fine-tuning each prompting method to enhance the model’s performance (Sec. 5.3). Lastly, we perform a speed analysis, comparing different prompting methods (Sec. 5.4).
All experiments use ground truth aggregated labels of the Assembly101-O dataset as the output of the predictor’s branches. This approach ensures that our performance analysis remains objective and untainted by errors from the step recognition branch, allowing for a more accurate evaluation of the step anticipator’s capabilities.
5.1 Model Selection
In this analysis, we evaluate the performance of various LLMs, as shown in Table 2 when using few-shot context. The results demonstrate a clear performance hierarchy among the tested models, with LLAMA 3.1 8B emerging as the superior performer across all metrics. LLAMA 3.1 8B achieved the highest accuracy at 47.40% and the best F1 score of 48.40%, significantly outperforming its counterparts. The performance trend correlates with model size and recency, suggesting that larger, more recently developed models benefit from architectural improvements and more extensive pre-training. This is exemplified by the LLAMA 3.1 8B model, which leverages its 8 billion parameters to capture more complex patterns and better generalize the assembly task. The superior performance of LLAMA 3.1 8B could be attributed to several factors, including potential advancements in its training methodology, the quality and diversity of its training data, and possible optimizations for general task comprehension. An exciting pattern observed across all models is the substantial disparity between Precision and Recall scores. The consistently high Recall scores, peaking at 91.30% for LLAMA 3.1 8B, indicate that these models identify relevant assembly steps. However, the markedly lower Precision scores suggest a tendency towards overprediction, resulting in false positives. This discrepancy highlights the inherent complexity of the assembly prediction task, where models must balance comprehensive step identification with selective precision.
Table 2: Results of different LLMs on Assembly101-O.
Model Accuracy Precision Recall F1
Mistral 7B 35.20 28.60 88.90 45.60
Gemma 9B 31.90 27.20 85.60 43.50
Phi 3 medium 4k instruct 39.60 29.50 88.70 46.20
LLAMA 2 7B 36.3 30.7 94.0 46.3
LLAMA 3 7B 36.40 29.80 87.30 45.60
LLAMA 3.1 8B 47.40 33.80 91.30 48.40
[…]
5.2 Prompt Analysis
System prompt.
The results in Table 3 highlight the positive impact of combining Automatic-Chain-of-Though (ACoT) reasoning with Few-Shot (FS) examples on the performance of LLAMA 3.1 8B across different modalities in the Assembly101-O dataset. In all prompting scenarios, the system prompt is written as:
{myframe}
System: I am going to provide an input sequence that represents a sequence of actions. Your task is to predict the next action of the last sequence based on the patterns observed in the provided input. Limit yourself to only answer with the predicted sequence, and follow the same format given as input.
This prompt consistently offers essential context about the task, improving the model’s performance, particularly in accuracy and precision.
In Zero-Shot (ZS) settings, the instruction prompt is presented as:
{myframe}
System: Below is an instruction that describes the task, paired with an input. Write a response that appropriately completes the request.
Few-Shot prompting enhances this by adding examples of input sequences and corresponding responses to guide the model.
{myframe}
Toy: ”bulldozer”
Input Sequence:
attach-cabin, attach-body, attach-track, attach-blade
Next Symbol:
attach-figurine
We explore two different input representations: a textual and a numerical one: in the first, we represent an action by its name and natural language, e.g., attach-wheel; in the second one, we use the index associated with the action triplet. For instance, in the textual modality, accuracy improves by 6.8%, and precision increases from 31.80% to 33.80%, reflecting a 6.3% gain. Similarly, in the numerical modality, including the system prompt boosts accuracy by 8.2% (from 40.10% to 43.40%) and precision by 4.6% (from 30.40% to 31.80%).
The improvements can be attributed to the system prompt and FS examples, which provide explicit context and structure, enabling the model to better understand the task’s nuances and generate more accurate and relevant predictions.
Table 3: Results of LLAMA 3.1 8B on Assembly101-O adding a system prompt.
System Prompt Modality Accuracy Precision Recall F1
Textual 47.40 33.80 91.30 48.40
Numerical 43.40 31.80 92.30 46.20
X Textual 44.40 31.80 87.30 46.10
Numerical 40.10 30.40 93.30 46.60
Prompting Methods.
When comparing ZS, FS, and ACoT prompting methods, ACoT emerges as the clear winner (Table 4), particularly when using the pre-trained model (Base), across both textual and numerical tasks. In the ACoT approach, automatic ACoT reasoning is employed, where the LLM is queried twice: first, to generate an internal reasoning step for the answer, and then to use this reasoning, combined with a few examples (as in FS), to derive the final answer. The first query is:
{myframe}
Let’s analyze each action in detail:
  1. 1.
    For the action ”attach the cabin,” consider the immediate effect and long-term consequences.
  2. 2.
    Repeat this analysis for each action in the sequence.
  3. 3.
    Identify patterns and causal relationships between the actions.
Now, let’s proceed with the analysis step-by-step:
This two-step process allows the model to provide an explanation for its decision and leverage the reasoning and examples to produce a more accurate and robust output. This superiority can be quantified: for example, in the textual modality, the ACoT base achieves a 51.40% accuracy, which is approximately 81% higher than the ZS base (28.40%) and 8.4% higher than the FS base (47.40%). In terms of the F1 score, the ACoT base (50.40%) outperforms the ZS base (42.40%) by 18.9% and the FS base (48.40%) by about 4.1%. This difference is less pronounced in numerical tasks, where the ACoT base achieves an accuracy of 40.40%, 47.4% higher than the ZS base (27.40%), and 8% higher than the FS base (37.40%).
These differences can be attributed to how these methods prompt the model. ZS and FS methods direct the model to generate answers straight from the input without guiding it through intermediate reasoning. While this may suffice for more straightforward tasks, it often falls short when dealing with more complex tasks that require deeper understanding.
ACoT, by contrast, excels in these scenarios because it prompts the model to break down the task into a series of logical, intermediate steps before concluding. This approach mirrors the “self-explanation” cognitive strategy, where breaking down and explaining each step improves problem-solving abilities. This explains why ACoT significantly outperforms ZS and FS, especially in tasks that demand nuanced understanding and intermediate reasoning.
Table 4: Results of LLAMA 3.1 8B on Assembly-101-O with different prompt methods, modalities, and inference modes.
Prompt Modality Mode Accuracy Precision Recall F1
ZS Textual Base 28.40 27.30 98.30 42.40
Textual Finetuned 28.40 27.30 99.30 42.40
Numerical Base 27.40 27.10 98.70 43.60
Numerical Finetuned 27.20 27.80 98.40 42.60
FS Textual Base 47.40 33.80 91.30 48.40
Textual Finetuned 46.40 30.80 97.30 45.40
Numerical Base 43.40 31.80 92.30 46.20
Numerical Finetuned 45.40 32.10 87.30 46.60
ACoT Textual Base 51.40 34.80 90.30 50.40
Textual Finetuned 46.40 31.80 85.30 47.60
Numerical Base 40.40 31.20 93.30 41.60
Numerical Finetuned 40.10 32.80 84.30 40.60
[…]
5.3 Finetuning Analysis
We analyze the performance of LLAMA 3.1 8B when fine-tuned on Assembly101-O (Table 4). Contrary to the common expectation that fine-tuning enhances performance by adapting the model to the task’s specific characteristics and distribution, this is not consistently observed. In both the ZS and FS methods, fine-tuning offers no significant improvement across metrics, with some cases showing similar or even lower precision and recall values. Notably, in the ACoT method, fine-tuning consistently results in a decrease in performance, particularly in terms of precision, suggesting that fine-tuning may not always provide the anticipated benefits.
This phenomenon can be explained by considering the dataset’s nature and ACoT prompting’s capabilities. Fine-tuning typically helps when there is sufficient data to prevent overfitting, allowing the model to generalize better. However, when the dataset is small, as in this scenario, fine-tuning may cause the model to overfit the specific examples seen during fine-tuning, capturing noise rather than functional general patterns.
For ACoT, the base model’s ability to handle complex reasoning through its intermediate steps seems sufficiently robust, and it does not benefit from the additional adjustments that fine-tuning provides. Fine-tuning might disturb these established reasoning pathways, leading to a decline in performance. This suggests that ACoT is inherently well-suited to tasks requiring deep reasoning, making it less dependent on the benefits of fine-tuning, especially in scenarios with limited data.
5.4 Speed Analysis
Table 5 presents a speed test comparison of different prompting methods—ZS, FS, ACoT—using LLAMA 3.1 on the Assembly101-O dataset. The results indicate that ZS is the fastest method, with an average processing speed of 0.208 seconds per sample. FS is slightly slower at 0.216 seconds per sample, representing a marginal 3.8% rise in processing time compared to ZS. ACoT, while offering superior accuracy and reasoning ability, is the slowest, taking 0.315 seconds per sample, which is a 51.4% increase in speed compared to ZS. These differences in speed can be attributed to the inherent complexity of each method. ZS and FS prompt the model to generate answers directly, with FS introducing a minimal overhead due to the additional context provided by the few examples. On the other hand, ACoT requires the model to engage in more complex, step-by-step reasoning, which naturally demands more computational resources and time. While ACoT’s performance benefits are clear, these come at the cost of slower processing, which could be a consideration in time-sensitive applications. Thus, the choice of prompting method should balance the trade-off between speed and output quality based on the task’s specific requirements.
Table 5: Speed Test with LLama 3.1 on Assembly-101-O
Model Speed (s/sample)
ZS 0.208
FS 0.216
ACoT 0.315
[…]
6 Aggregation Strategy
6.1 Step Recognizer Analysis
The action recognizer is the architecture module responsible for online action detection, a task extensively explored in recent research due to its various possible applications. Several models have been proposed using datasets such as Thumos [52], TVSeries [53], and Fine Action [54], with the current leading model being MiniROAD [32]. We rely on the state-of-the-art MiniROAD, even though it should be considered that its performances on procedural datasets has been suboptimal, largely due to two key factors: the shift in action distribution between the training and testing phases, and the inherent class imbalance within the dataset. The distribution shift occurs because the actions seen during training often differ from those encountered during testing, leading to poorer generalization. Additionally, Assembly101 contains a significant imbalance in class frequencies (see Figure 4), where common actions dominate the dataset while less frequent yet crucial procedural actions are underrepresented.
Figure 4: 20 most frequent classes in Assembly101-O
6.2 Frame Aggregation
While MiniROAD [32] produces per-frame predictions, the anticipator expects a sequence of distinct actions as input. A standard approach for transforming per-frame predictions into action sequences involves aggregating identical neighboring predictions into a single action segment. However, MiniROAD’s performance issues introduce significant noise into the input sequence for the anticipator module, mainly due to inconsistencies in per-frame predictions, such as oscillating between similar actions, which leads to unstable action predictions and degrades results in subsequent stages.
To address this, we propose different frame aggregation strategies to efficiently convert per-frame predictions into coherent action sequences, aiming to reduce noise and stabilize predictions. This results in a cleaner input for the anticipator module. Our approach introduces three distinct aggregation methods, as shown in Figure 5.
Figure 5: Predictions bars: (a) Ground truth predictions, (b) Predictions using the recognizer,(c) First aggregation strategy, (d) Second aggregation strategy, (e) Third aggregation strategy
Each aggregation strategy’s performance will be evaluated using the Levenshtein Similarity between the predicted and ground truth action sequence, which derives from the Levenshtein distance. The similarity between two action sequences A and B is defined as:
Levenshtein $\text{Levenshtein Similarity}(A,B)=1-\frac{d_{dev}(A,B)}{\max(len(A),len(B))}$ (4)
where: $d_{dev}(A,B)$ is the Levenshtein distance between procedure A and B, which is the minimum number of single-action insertions, deletions, or substitutions required to transform procedure A into procedure B, and len(A) and len(B) represent the length of procedure A and B, respectively. Note that the term max(len(A),len(B)) normalizes the score by the length of the longer procedure, ensuring that the ratio is between 0 and 1.
6.2.1 First strategy
The first aggregation strategy (Figure 5.c) we propose consists of a two-step process. In the first step, we calculate the mode for each non-overlapping sliding window of frames of a given length $l$. Because the windows are non-overlapping, this approach introduces a slight delay in processing, which is a trade-off for improved prediction stability. The mode, representing the most frequent action label within that window, replaces all predictions in that window, ensuring consistency across the frames and removing the noise of smaller portions wrongly predicted. In the second step, we apply the elimination of successive duplicates to remove consecutive repeated action labels, which prevents redundant predictions. This cleaned output serves as the input to the anticipation module, ultimately aiming to produce a more stable and accurate sequence of action predictions.
6.2.2 Second strategy
The second aggregation strategy (Fig. 5.d) we propose employs a sliding window approach with a stride of $1$ frame, allowing for overlapping portions of frames. In this method, we evaluate the mode within the overlapping window that includes the last frame and the previous $l-1$ frames. The prediction of the last frame is then substituted with this mode, which captures the most frequent action label in that localized area. This method is designed with the intent to reduce delay, as it processes frames more dynamically compared to non-overlapping windows.
6.2.3 Third Strategy
The third aggregation strategy (Fig. 5.e) uses a stride of 1 frame, similar to the previous approach. To address the potential issue of the mode being incorrectly positioned within the window, we opt to substitute the prediction of the central frame rather than the final frame, with some corrections for the initial and final frames in the video. This adjustment ensures that the most frequent action label is applied to a frame that is more representative of the overall context, thus enhancing the accuracy of the prediction.
Figure 6: Levenshtein similarity for the proposed aggregation strategies. The maximum similarity is achieved by the first strategy when using a window size of 500 frames.
Figure 7: Histogram of the durations of the action in Assembly101-O
Figure 6 illustrates that the first approach is the most effective in maximizing similarity.
Considering that datasets like Assembly101 are recorded at 30 fps and the average action lasts about 570 frames (Fig. 7), we selected 500 frames as a sensible trade-off between achieving high similarity and minimizing the computational burden.
7 Experiments
7.1 Baselines
To estimate the effectiveness of our model, we evaluate its performance by comparing it against several baseline models based on the metrics presented in Sec. 4.1. This evaluation framework builds upon the baselines initially introduced in PREGO [22], as detailed below:
One-step memory In PREGO [22], we defined a transition matrix considering only the correct procedures. Specifically, given the set of actions $\mathcal{A}$ in the training set with $|\mathcal{A}|=C$, we define a transition matrix $M\in\mathbb{R}^{C\times C}$, which records, in position $(l,m)$, the number of occurrences where action $m$ follows action $l$. During evaluation, an action in the test split that does not correspond to a transition recorded in this training set matrix is labeled as a mistake. This baseline provides a simple method for recognizing deviations from standard procedural transitions.
OadTR for mistake detection The work in [14] presents a framework for online action detection called OadTR, which uses a Vision Transformer to capture video clips’ temporal structure and context. The framework consists of an encoder-decoder architecture: the encoder processes historical observations to output a task token that represents the current action, while the decoder refines this task token by incorporating information from anticipated future clips. In the context of procedural error detection, we consider a mistake to have occurred if the output from the encoder does not match the output from the decoder, indicating a divergence between observed and expected actions.
BERT-based Mistake Detection We also leverage the capabilities of BERT [55], using its specific [CLS] token to predict whether an action sequence is correct or erroneous. More specifically, we fine-tune BERT with a next-sentence-prediction task, training the model to determine whether step B logically follows step A within a given procedure. In our setting, each step is represented as a set of two words, such as attach wheel, which describes coarse-level actions. During evaluation, BERT is presented with pairs of actions and predicts whether the sequence of those actions is correct. The advantage of BERT lies in its pre-training on a vast corpus of text, which helps it understand contextual relationships between procedural steps, making it effective for classifying sequences and understanding the logical flow of actions.
Learned Task Graph Representations We also evaluate an approach based on explicit task graph representations [16]. Procedural activities consist of sequences of key steps that work toward specific goals, making task graphs a valuable, human-understandable representation for capturing the partial ordering of these steps. Unlike traditional methods that rely on hand-crafted procedures to extract task graphs from videos, this approach uses maximum likelihood optimization of edge weights to directly learn task graphs, allowing gradient-based learning that can integrate seamlessly with neural networks. While effective, this baseline is not directly comparable to our method, as it requires access to the entire dataset during training to extract the task graph, whereas our approach relies only on few-shot inference for anticipating actions. We refer to the “Direct Optimization” method trained with the “Task-Graph Maximul Likelihood” loss proposed in [16], which we refer to as TGML-DO.
In addition to these baselines, we introduce several new baselines to explore corner-case scenarios, offering a broader perspective for comparison:
Best-Case Scenario In the best-case scenario, we assume that the anticipation branch always predicts the correct action. When the procedural context indicates a mistake is expected, the model will predict anything other than the incorrect action, effectively identifying the mistake. Conversely, for correct actions, the model will consistently make accurate predictions. This ensures that all mistakes are flagged correctly and all correct actions are anticipated without error, serving as an upper bound for the system’s performance.
Worst-Case Scenario In the worst-case scenario, the anticipation branch always predicts the incorrect action. Specifically, if a mistake is expected, the model fails to detect it, and if a correct action is expected, the model still predicts an incorrect action. This serves as a lower bound, helping us understand the limitations of our approach under adversarial conditions.
Random Anticipation In the random scenario, the anticipation branch provides random guesses for each action, regardless of the procedural context. This baseline highlights the importance of a well-informed anticipation module by comparing it against a model without contextual awareness of the sequence.
These baselines allow us to comprehensively evaluate our model’s performance across a range of conditions, from ideal to highly challenging scenarios, providing a fair and robust comparison to gauge the effectiveness of our proposed approach.
7.2 Sequence-Level Results
We evaluate our model, which utilizes Automatic Chain of Thought for intermediate reasoning, alongside various baseline methods on two datasets, Assembly101-O and Epic-tent-O, as shown in Table 2. For this evaluation, we replaced the step recognition predictions with ground truth action labels (Oracle setting) to establish an upper bound on performance, simulating perfect action recognition by the video branch.
The comparison includes several baselines. The One-step memory method only considers the previous action, while BERT takes a more abstract reasoning approach, leveraging past actions more effectively. However, this leads to a conservative bias, reducing false alarms and missing some mistakes. PREGO leverages symbolic reasoning to model richer contextual information, with PREGOLlama demonstrating the strongest performance across most metrics.
In the Assembly101-O dataset, our model achieves a balanced performance, particularly in the F1-score, outperforming many baseline methods. Similarly, on the Epic-tent-O dataset, our model attains competitive precision and recall, highlighting its effectiveness in detecting procedural mistakes across different tasks. Specifically, our model outperforms OadTR, which suffers from fixed video segment processing limitations and benefits from leveraging semantic labels directly in reasoning processes, resulting in a 102%
The TGML-DO method [16], which employs task graph representations learned through gradient-based optimization, also shows competitive performance. However, its reliance on the entire dataset during training contrasts with our approach, which relies on few-shot inference. This difference may contribute to our model’s superior performance when using predictions for the action recognizer in the Assembly101-O dataset, as our method is more robust to procedural variations and better at generalizing to unseen cases.
The results indicate that incorporating the Automatic Chain of Thought enhances intermediate reasoning, allowing our model to anticipate actions more effectively and detect deviations from normal procedure patterns. This positions our model as a powerful alternative, with better procedural mistake detection performance than symbolic reasoning-based approaches like PREGO and other baselines.
Table 6: A comparative assessment between Ours and the chosen baseline methods is conducted to detect procedural mistakes using the Assembly101-O and Epic-tent-O datasets.
Assembly101-O Epic-tent-O
Step Recog. Step Antic. Precision Recall F1 score Precision Recall F1 score
Random Oracle 45 91 25 8.4 100 0.16
One-step memory Oracle 16.3 30.7 21.3 6.6 26.6 10.6
BERT [55] Oracle 78.2 20.0 31.8 75.0 5.6 10.4
PREGO [22] Oracle Llama 2 30.7 94.0 46.3 10.7 86.7 19.1
TI-PREGO Oracle Llama 3.1 34.80 90.30 50.40 10.6 93.5 22.4
TGML-DO [16] Oracle 46.7 90.40 61.60 20 27.3 23.1
TGML-DO [16] MiniRoad [32] 22.50 37.30 28.10 73.3 14.1 23.7
OadTR for MD [14] OadTR [14] OadTR [14] 24.3 18.1 20.7 6.7 21.7 10.2
PREGO [22] OadTR [14] Llama 2 22.1 94.2 35.8 9.5 93.3 17.2
PREGO [22] MiniRoad [32] Llama 2 27.8 84.1 41.8 8.6 20.0 12.0
TI-PREGO MiniRoad [32] Llama 3.1 30.30 76.8 43 9.8 100 19.2
7.3 Frame-Level Results
Figure 8: With a window frame of size $n>1$ there can be a mismatch between the ground truth window, whose distribution is often not constant, and the expanded prediction, whose is.
In Section 6.2, we explained how the output from the step recognizer is aggregated before being passed to the anticipator. In this section, we evaluate the anticipator’s output at the frame level, matching the frame-level input the step recognizer receives.
During aggregation, action sequences are split into windows of fixed size $n$, where the most frequent action within each window is selected as the representative action. To further minimize repetitions, these aggregated actions are refined to discard duplicates, resulting in a sequence of unique actions, as depicted in the top part of Fig. 8.
We then compute the previously defined metrics at the frame level. To reverse the aggregation process, we expand the anticipator’s output back to individual frames. This is achieved by keeping track of the original number of frames within each window during the aggregation step. However, due to this expansion, when aggregation is performed (i.e. $n>1$), even the best-case scenario does not necessarily achieve an optimal performance of an F1 score of $1$, nor does the worst-case scenario result in an F1 score of 00. This inconsistency between the expanded predictions and ground truth sequences is illustrated in Fig. 8.
To provide a more comprehensive evaluation, we include balanced accuracy (see Eq. 3) in Table 7, where we investigate the effect of different window sizes on the Assembly101-O dataset. For clarity, we report the results for corner cases alongside our model’s performance, allowing for a more contextual understanding of the outcomes. These corner cases are further detailed in Sec. 7.1. We also report the results for corner cases alongside our model’s performance, allowing for a more contextual understanding of the outcomes. These corner cases are further detailed in Sec. 7.1.
Table 7: Performance metrics for varying frame window sizes on the Assembly101-O dataset. We focus on giving an upper bound, so the anticipator always takes grounds truth data as an input. We report balanced metrics to ensure a fair comparison, given the significant imbalance in accuracy and precision.
Frame Window Accuracy Precision Recall F1 Balanced Accuracy
Best 1 100.00 100.00 100.00 100.00 100.0
Worst 1 0.00 0.00 0.00 0.00 0.0
Random 1 0.25 0.05 100.00 0.10 50.1
Ours 1 51.50 0.06 62.64 0.12 57.1
Best 50 98.42 2.95 100.00 5.73 99.2
Worst 50 0.00 0.00 0.00 0.00 0.0
Random 50 0.35 0.05 100.00 0.10 50.1
Ours 50 83.93 0.08 25.27 0.15 54.6
Best 200 93.93 0.78 100.00 1.55 97.0
Worst 200 0.01 0.00 0.00 0.00 0.0
Random 200 0.44 0.05 98.35 0.09 49.4
Ours 200 83.07 0.06 20.33 0.11 51.7
Best 400 88.08 0.40 100.00 0.80 94.0
Worst 400 0.03 0.00 0.00 0.00 0.0
Random 400 0.40 0.05 100.00 0.10 50.2
Ours 400 83.87 0.06 21.43 0.13 52.7
All predictions are made online, similar to the sequence-level approach. The key difference is that in the frame-level scenario, inputs are not aggregated, and the anticipator operates at each frame independently, predicting the action for each frame. In scenarios where aggregation is used, the resulting inconsistencies introduce challenges, as discussed earlier. For instance, with a window size of $200$, and considering an idealized (oracular) anticipator, we observe $0.05\%$ true positives, $6\%$ false positives, $94\%$ true negatives, and 00 false negatives across a total of 380,110 frames.
8 Limitations
Our dual-branch architecture for online procedural mistake detection demonstrates promising results. In this section, we acknowledge some of its limitations.
Real-time processing constraints can present challenges, particularly when incorporating Large Language Models (LLMs), which may introduce latency during inference. Our focus is maintaining accuracy in an online setting, rather than optimizing for real-time performance. Nonetheless, real-time optimization remains a promising direction for future work.
While symbolic reasoning simplifies certain aspects of action prediction, it may not fully capture the complex semantic relationships between actions. Future research could explore hybrid approaches that integrate symbolic reasoning with richer representations.
Lastly, our evaluation metrics primarily emphasize action recognition and anticipation accuracy. However, other aspects, such as robustness to noise, the ability to handle ambiguous scenarios, and the interpretability of predictions, could also be valuable considerations. Addressing these limitations would help further advance procedural mistake detection and improve its effectiveness in dynamic and diverse environments.
9 Conclusion
This paper explored the challenging task of detecting procedural errors in egocentric videos online, focusing on a dual-branch architecture that combined action recognition and anticipation. By leveraging Large Language Models (LLMs) within the action anticipation branch, we used in-context learning and Automatic-Chain-of-Thought to demonstrate their ability to predict future steps based on prior actions. Our experiments comprehensively analyzed various prompting schemes and aggregation strategies to optimize performance in online action recognition, highlighting the importance of effective prompt formulation for LLM-based anticipators.
Our results emphasized the challenges of per-frame evaluation in real-time action recognition systems and showed how symbolic reasoning, combined with predictive models, enhanced procedural mistake detection.
Through extensive experimentation, we demonstrated the effectiveness of our dual-branch architecture, achieving new state-of-the-art performance for online mistake detection. This underscored the potential of integrating action recognition and anticipation in a unified framework for detecting procedural errors as they unfolded online. Future work could explore further refinements in symbolic representation and expand the range of procedural tasks and environments in which our framework can be applied.
Acknowledgements
This work was carried out while Leonardo Plini was enrolled in the Italian National Doctorate on Artificial Intelligence run by Sapienza University of Rome. We thank DsTech S.r.l. and the PNRR MUR project PE0000013 Future Artificial Intelligence Research (FAIR) (CUP: B53C22003980006 and CUP: E63C22001940006) for partially funding the Sapienza University of Rome and University of Catania.
Data Availability Statement
The datasets generated or analysed during the current study are available in (1) https://sites.google.com/view/epic-tent and (2) https://assembly-101.github.io/